1, You want to save to buy a laptop that costs $1,122. Therefore, you deposit $673 into a bank that pays 14.28% interest compounded annually. How many year will it take for you to buy the laptop in your shopping list?

Please round your answer to the second decimal, e.g. 1.50 (year).

2,

If you deposit $909 in a bank account today that pays an annual interest rate of 4.41% compounded monthly, how much will you expect to have in your bank account after 3 years?

Please round your answer to the second decimal without dollar sign (e.g. 0.00)
Answer #1

1.We use the formula:  
A=P(1+r/100)^n
where
A=future value
P=present value  
r=rate of interest
n=time period.

1,122=673*(1.1428)^n

(1,122/673)=(1.1428)^n

Taking log on both sides;

log (1,122/673)=n*log (1.1428)

n=log (1,122/673)/log (1.1428)

=3.83 years(Approx)

2.We use the formula:  
A=P(1+r/12)^12n
where
A=future value
P=present value  
r=rate of interest
n=time period.

A=909*(1+0.0441/12)^(12*3)

=909*1.14117393

=1037.33(Approx)